# Large Language Model Performance Benchmarking on Mobile Platforms: A Thorough Evaluation

### v1

@@ Abstract
As large language models (LLMs) increasingly integrate into every aspect of our work and daily lives, there are growing concerns about user privacy, which push the trend toward local deployment of these models. There are a number of lightweight LLMs (e.g., Gemini Nano, LLAMA2 7B) that can run locally on smartphones, providing users with greater control over their personal data. As a rapidly emerging application, we are concerned about their performance on commercial-off-the-shelf mobile devices. To fully understand the current landscape of LLM deployment on mobile platforms, we conduct a comprehensive measurement study on mobile devices. We evaluate both metrics that affect user experience, including token throughput, latency, and battery consumption, as well as factors critical to developers, such as resource utilization, DVFS strategies, and inference engines. In addition, we provide a detailed analysis of how these hardware capabilities and system dynamics affect on-device LLM performance, which may help developers identify and address bottlenecks for mobile LLM applications. We also provide comprehensive comparisons across the mobile system-on-chips (SoCs) from major vendors, highlighting their performance differences in handling LLM workloads. We hope that this study can provide insights for both the development of on-device LLMs and the design for future mobile system architecture.
1 Introduction
@@
As a rapidly emerging application, LLMs differ significantly from traditional mobile applications in terms of computational complexity and memory demands. Therefore, we are particularly concerned with their performance on commercial off-the-shelf (COTS) mobile devices, as well as the performance fluctuations caused by mobile operating systems and hardware architecture compatibility. Several studies have taken the first steps to explore the current state of mobile LLM[25][27][34]. They have conducted preliminary assessments of model deployment and inference performance through the development of specialized benchmark suites. However, they mainly focus on performance metrics such as token throughput and latency, without looking into the underlying hardware dynamics and system configurations, such as CPU/GPU utilization, number of threads, operating frequencies, etc. Besides, there is a lack of analysis about how these hardware and system-level factors affect on-device LLM inference. This is essential for developers to identify and address bottleneck for mobile LLM applications.
@@
In this paper, we aim to reveal the current state and potential of mobile hardware for supporting on-device LLMs by deploying LLMs across a diverse range of mobile devices. We select multiple high-tier SoCs from different vendors that represent currently available resources and development trends in mobile hardware. We analyze the differences in SoCs from major vendors and ultimately select SoCs represented by Qualcomm, HiSilicon, and MediaTek to perform local LLM inference. We deploy a B model on mobile devices with llama.cpp [14] and MLC LLM [38], which are the two popular mobile LLM inference engines. During inference, we collect comprehensive metrics with specific profilers including Snapdragon Profiler [35] and Arm Streamline [5] to make sure that all the data accurately reflect the hardware performance. We investigate how the hardware specifications, including big.LITTLE cores[6], cache size, and memory bandwidth, affect the inference performance, which is essential to help the development of future mobile SoCs to accelerate LLMs.
@@
From our study, we have the following key observations:
1.
While LLMs are typically considered to be resource-intensive, the current implementation does not fully explore the potential of mobile processors. It only utilizes of the arithmetic units on mobile GPUs.
2.
Processors with superior hardware specifications may not necessarily exhibit better performance (e.g., Mali-G720 Immortalis MP12 vs Adreno 750). In addition, the Adreno GPUs consistently outperform the Mali GPUs in overall performance.
3.
@@
Given the "big.LITTLE" architecture of mobile CPUs, in most cases, setting the number of threads equal to the number of big cores (prime and performance cores) can achieve the optimal performance. Adding more cores (efficiency cores) may degrade the performance.
4.
@@
Specialized machine instructions, such as smmla and sdot in Arm architecture, significantly enhance LLM performance. By carefully rearranging the weight matrix, these instructions can yield up to a speed improvement.
5.
Specialized AI accelerators, such as the Hexagon NPU, demonstrate a remarkable prefill speed, representing a -fold increase compared to CPU and GPU. However, their performance in decoding does not exhibit a similar level of enhancement.
@@
The rest of the paper is organized as follows. In Section 2, we present preliminaries and related works on this topic. In Section 3, we describe our measurement setup and methodology. In Section 4, we present performance metrics that are of primary concern to users, while Section 5 discusses findings related to the underlying hardware characteristics and inference framework, which may interest developers. We discuss potential improvements in Section 6 and explain the remaining issues in Section 7. Finally, we conclude the paper in Section 8.
2 Related Works
2.1 Light-Weight LLMs
@@
2.2 On-device Deployment of LLMs
@@
There are some inference frameworks widely used including llama.cpp [14], MLC LLM [38], mnn-llm [3], mllm [44], etc. Among them, llama.cpp and mnn-llm are optimized for LLM inference on CPUs, while MLC LLM leverages the powerful computational capabilities of GPUs. Despite the availability of dedicated AI accelerators such as NPUs, APUs, and TPUs in the latest generation of mobile SoCs [20], developing for these accelerators presents challenges due to the typically closed-source nature of vendor-specific SDKs. Currently, only mllm and PowerInfer-2 [46] claim to support Qualcomm NPUs for accelerating LLM inference on mobile devices. The development of LLM solutions for AI accelerators with dedicated architectures is still an ongoing challenge for developers.
In addition to those open-source LLM inference engines, mobile SoC vendors and manufacturers have also realized the importance of on-device LLM inference. Qualcomm, the vendor of Snapdragon SoCs, asserts that it can accelerate Llama-2 7B and Llama-3 8B models using the NPU on Snapdragon Gen2 and Gen3 platforms [36]. Similarly, MediaTek has announced optimizations for Llama 2 Generative AI applications using its APU on the Dimensity 9300 and 8300 [31]. However, both solutions remain closed-source and have yet to be validated through real-world applications.
2.3 Existing Measurement Studies
@@
However, none of these studies provide in-depth analysis about the impact of mobile hardware during inference, such as accelerator utilization, memory bandwidth, and fluctuations in CPU frequency. Although [27] touches on the impact of SoCs, it only examines the macro differences of SoCs from a single vendor. In contrast, our approach involves testing several SoCs from multiple vendors to gain a comprehensive understanding of how mainstream mobile processors support LLMs. This will help us identify hardware bottlenecks and potential future upgrades necessary for optimizing LLM performance on mobile devices.
@@
XiaoMi Pad6 Pro | Snapdragon 8+ Gen1 | Cortex-X2/710/510 | Adreno 730 | 12GB/16GB | top-tier | Android13 | 2022.5 |
Huawei Matepad11 Pro | Snapdragon 870 | Cortex-A77/A77/A55 | Adreno 650 | 5.5GB/8GB | mid-tier | Harmony 4 | 2021.1 |
@@
Huawei Nova7 | Kirin 985 | Cortex-A76/A76/A55 | Mali-G77 MP8 | 4GB/8GB | mid-tier | Harmony 4 | 2020.4 |
3 Experimental Setup and Methodology
3.1 Experimental Setup
Testing Devices. Table 1 shows the devices used in our measurement. There are mobile devices from three SoC vendors: Snapdragon, MediaTek, and Hisillicon. All CPUs feature eight Arm Cortex cores but differ in implementation details such as big.LITTLE configurations and cache sizes. It is noteworthy that the Snapdragon 8 Gen3 and Dimensity 9300 represent the most advanced SoCs available for mobile devices. We also test devices ranging from high-tier to mid-tier to capture a broad spectrum of hardware heterogeneity. In terms of GPU, the vast majority of mobile GPUs are from Adreno (Qualcomm) and Mali (Arm). Thus, we designate Snapdragon SoCs as representatives of Adreno GPUs and the rest as representatives of Mali GPUs.
@@ 3.1 Experimental Setup
Frameworks and Engines. In this paper, we would like to reveal the performance of LLM on a variety of different accelerators. To achieve this, we select two representative frameworks: llama.cpp for CPU deployments and MLC LLM for GPU deployments, to assess the inference capabilities of general accelerators. For novel AI accelerators (e.g., NPU), rather than conducting our own device deployments, we refer to existing studies for results of mllm and PowerInfer-, which are only available for Qualcomm NPUs. These references will be discussed in subsequent sections.
@@
MLC LLM [38] is a high-performance LLM deployment engine built on the TVM machine learning compiler. Unlike llama.cpp, which focuses on optimizing for mobile CPUs, MLC LLM is designed to harness the power of mobile GPUs. Leveraging TVM, MLC LLM can automatically generate and optimize operators for a specific model and backend. These operators are then loaded and executed via the TVM Runtime. For Android-based GPUs, MLC LLM utilizes the OpenCL backend to ensure efficient execution.
@@
Measurement toolkits. In order to collect fine-grained hardware metrics such as real-time CPU frequency and memory bandwidth utilization, we choose Perfetto [1], Snapdragon Profiler [35] and Arm streamline [5] as our primary tools for monitoring hardware dynamics during inference. Perfetto offers comprehensive system-wide performance traces from Android devices, including data from the kernel scheduler, userspace instrumentation, and other sources. For CPU monitoring, Perfetto captures frequency, utilization, and event scheduling for each CPU core. Snapdragon Profiler and Arm Streamline are vendor-specific profilers used to track GPU behavior, including arithmetic and load/store unit utilization. Specifically, Snapdragon Profiler provides detailed timelines of OpenCL kernel executions, which is valuable for operator-level analysis.
Evaluation Metrics. We collect results from two main aspects. First, we gather performance metrics that are of interest to users, including prefill and decoding speed, memory footprints, and battery consumption. These results are present in Section 4. Second, we analyze hardware dynamics, focusing on real-time processor utilization and working frequency, impact of Dynamic Voltage and Frequency Scaling (DVFS) and operator implementation, which developers are concerned about. These results are present in Section 5. They provide clues for the performance metrics and can help us identify potential hardware and operating system bottlenecks during inference.
3.2 Workflow and Methodology
@@
For llama.cpp, we build the model library and compile the source code into an executable file on each device individually, using CMake and the Android NDK. All the compile options are kept the same by default. When necessary, we adjust the compile options to enable specialized instructions to ensure optimal performance (Section 5.1.2). Since it relies on basic C libraries, no additional environment setup is required. After transferring model weights in GGUF format to the device, inference can be performed easily using the ADB command-line tool. For MLC LLM, there is a native application which TVM runtime and necessary libraries are packed in. The APK can be installed on the device, allowing interaction with the LLM through a graphical interface. MLC LLM cross-compiles the LLM models for the mobile platform, and on all devices, the runtime version including tvm and java is the same.
@@
4 Performance: Users’ Perspectives
@@
4.1 Token Throughput and Memory Footprints
4.1.1 Performance on CPUs
@@
Figure 3 shows the performance of CPUs, including prefill speed and decoding speed with llama.cpp. Overall, there has been a consistent improvement in both prefill and decoding speeds with new-generation processors. The Dimensity on Vivo Pad3 Pro, being the most advanced new-generation SoC, demonstrates the highest performance, achieving over speedup in prefill and nearly speedup in decoding compared to the older SoC Snapdrgon on the Huawei Matepad Pro. The Snapdragon 8 Gen (XiaoMi Pro) ranks second among all tested devices, showing % throughput of the Dimensity . Comparing Snapdragon SoCs across different tiers, each successive generation shows approximately a improvement in prefill speed, while decode speed improves even more significantly, ranging from to . We note that Snapdragon 870 (on Huawei Matepad11 Pro) and Kirin 9000E (on Huawei Matepad12.6 Pro) consist of the same cores, and the performance gap may result from higher CPU frequency on Kirin 9000E. Notably, a significant performance gap persists between Kirin SoCs and those from Snapdragon and MediaTek in terms of supporting LLMs.
@@
Our results highlight a notable performance disparity in LLM inference across mobile CPUs. Specifically, top-tier SoCs like the Dimensity demonstrate significantly better LLM support compared to mid-tier SoCs. This suggests that for optimal LLM performance, developers should focus on leveraging the most advanced processors available. Furthermore, top-tier SoCs support advanced machine instructions that can significantly boost performance. We will examine this potential in detail in Section 5.1.2.
@@
We further evaluate the performance with two scenarios: one involving a prompt of tokens with a generation of tokens, and the other involving a prompt of tokens with a generation of tokens. The larger input matrix in the latter case implies greater computational intensity and resource demands. Figure 3 illustrates that inference speed decreases as the token length increases. The results may be attributed to the impact of DVFS. With longer prompts and generation tasks, the increased load causes the CPU frequency to throttle more aggressively. Across all tested devices, the performance degradation for both prefill and decoding is relatively consistent, ranging between and .
@@
In addition to latency, memory footprint is a also crucial metric. On all six devices, memory usage remains consistently around GB and shows minimal variation during inference once the model weights are loaded into main memory from external storage (e.g., Universal Flash Storage, UFS). This consistency is due to the fact that llama.cpp applies for a fixed-size memory block during initialization, which remains constant once allocation is complete. The allocated memory is then segmented into various components for different variables, including model weights, KV-cache, buffers, context and etc.
4.1.2 Performance on GPUs

### v3

# Understanding Large Language Models in Your Pockets: Performance Study on COTS Mobile Devices

@@ Abstract
As large language models (LLMs) increasingly integrate into every aspect of our work and daily lives, there are growing concerns about user privacy, which push the trend toward local deployment of these models. There are a number of lightweight LLMs (e.g., Gemini Nano, LLAMA2 7B) that can run locally on smartphones, providing users with greater control over their personal data. As a rapidly emerging application, we are concerned about their performance on commercial-off-the-shelf mobile devices. To fully understand the current landscape of LLM deployment on mobile platforms, we conduct a comprehensive measurement study on mobile devices. While user experience is the primary concern for end-users, developers focus more on the underlying implementations. Therefore, we evaluate both user-centric metrics—such as token throughput, latency, and battery consumption—and developer-critical factors, including resource utilization, OS strategies, and inference engines. We also provide comprehensive comparisons across the mobile system-on-chips (SoCs) from major vendors, highlighting their performance differences in handling LLM workloads, which may help developers identify and address bottlenecks for mobile LLM applications. We hope that this study can provide insights for both the development of on-device LLMs and the design for future mobile system architecture.
Index Terms:
Large Language Models, Measurement, On deviceI Introduction
@@
As a rapidly emerging application, LLMs differ significantly from traditional mobile applications in terms of computational complexity and memory demands. Therefore, we are particularly concerned with their performance on commercial off-the-shelf (COTS) mobile devices, as well as the performance fluctuations caused by mobile operating systems and hardware architecture compatibility. Several studies have taken the first steps to explore the current state of mobile LLM[8][9][10]. They have conducted preliminary assessments of model deployment and inference performance through the development of specialized benchmark suites. However, they mainly focus on performance metrics such as token throughput and latency, without looking into the underlying hardware dynamics and system configurations, such as CPU/GPU utilization, operating frequencies, etc. Besides, there is a lack of analysis about how these hardware and system-level factors affect on-device LLM inference. This is essential for developers to identify and address bottleneck for mobile LLM applications.
@@
In this paper, we aim to reveal the current state and potential of mobile hardware for supporting on-device LLMs by deploying LLMs across a diverse range of mobile devices. We select multiple high-tier SoCs from different vendors that represent currently available resources and development trends in mobile hardware. We analyze the differences in SoCs from major vendors and ultimately select SoCs represented by Qualcomm, HiSilicon, MediaTek and Apple to perform local LLM inference. We deploy a B model with -bit quantization on mobile devices with llama.cpp [11] and MLC LLM [12], which are the two popular mobile LLM inference engines. During inference, we collect comprehensive metrics with specific profilers including Snapdragon Profiler [13], Arm Streamline [14] and Xcode [15] to make sure that all the data accurately reflect the hardware performance. We investigate how the hardware specifications, including big.LITTLE cores[16] and memory bandwidth, affect the inference performance, which is essential to help the development of future mobile SoCs to accelerate LLMs.
Considering the diversity of mobile hardware and rapid iteration of capability, current research exhibits a notable gap in both comprehensive understanding and detailed evaluation of LLM deployment on mobile devices. In order to show the usability of on-device inference and provide developers with valuable insights, we have summarized the following key questions:
Q1. Hardware Requirements: What are the key hardware requirements (e.g., computation capability, RAM size) to support LLMs on mobile devices?
Q2. Resource Configuration: What is the optimal runtime resource configuration, including system policies (DVFS and thermal throttling), the number of parallel threads, core affinity settings, and others? Moreover, how does the optimal configuration differ when inference runs exclusively versus concurrently with other tasks? Furthermore, how can this configuration be adapted to different devices?
Q3. Software-Hardware Alignment: Do the implementations and optimizations of inference frameworks align with general-purpose processors (CPUs) or specialized accelerators (GPUs) from various vendors? How much do software implementations limit the utilization of hardware capabilities, such as computation throughput or memory bandwidth?
@@
From our study, we have the summarized corresponding observations to the above questions:
Response to Q1:
CPU performance: The iOS device shows a clear advantage, particularly in the prefill stage. In the decode stage, while the Apple M1 reaches about 10 tokens/s, non-iOS devices with Armv8-A CPUs (e.g., Cortex-A76 and A77) only achieve 2–4 tokens/s. In contrast, Armv9-A CPUs (e.g., Cortex-X4), equipped with specialized instructions such as smmla and sdot, deliver substantial gains and can reach throughput comparable to iOS devices.
Performance of hardware accelerators: As for accelerators, Specialized AI accelerators, such as the Hexagon NPU, demonstrate remarkable prefill speed, achieving a -fold increase compared to CPU and GPU on the same SoC. For decoding on GPU, due to the memory-bound nature of the decode stage, accelerating generation remains challenging.
RAM requirement: The memory (RAM) requirement to support LLMs depends on the model size and the quantization bit-width. A -bit quantized B-parameter model typically requires at least GB of RAM to store the model weights and maintain a minimal context (eg. KV cache).
Battery support: For inference on modern high-tier mobile devices (e.g., Huawei Matepad Pro), the power consumption is around mAh per round. Typical smartphone batteries have a capacity of mAh to mAh, they can support hundreds of rounds of inference.
Response to Q2:
DVFS Strategy: The DVFS strategy could be more aggressive. As inference on mobile devices is typically intermittent, it is acceptable to increase the CPU frequency to speed up for single-round inference, as users are particularly sensitive to the first token generation latency. In contrast, GPUs on non-Apple devices tend to operate at a relatively fixed frequency (not the maximum). Therefore, there is still room for further acceleration when performing inference on GPUs.
@@
Computing Thread: Given the “big.LITTLE” architecture of mobile CPUs, in most cases, setting the number of threads equal to the number of big cores (prime and performance cores) can achieve the optimal performance. Adding more cores (efficiency cores) may degrade the performance. When running in parallel with other tasks, the optimal configuration should vary dynamically, depending on both the real-time workload size and the nature of the co-running tasks.
Response to Q3:
@@
Instruction speedup: Specialized machine instructions, such as smmla and sdot in Arm architecture, can significantly improve the performance. By carefully rearranging the weight matrix, these instructions can yield up to a 4× speed improvement.
Low GPU utilization on non-Apple SoC: While LLMs are typically considered to be resource-intensive, the current implementation does not fully explore the potential of most mobile processors. It only utilizes % % of the arithmetic units on mobile GPUs in the prefill stage.
Gap between Adreno and Mali GPUs: Due to unaligned software implementations, processors with superior hardware specifications may not necessarily exhibit better performance (e.g., Mali-G Immortalis MP vs Adreno ). In addition, the Adreno GPUs consistently outperform the Mali GPUs in overall performance.
Memory bottleneck: Due to the memory-bound nature of decoding, the decode speed does not vary significantly across top-tier devices. In contrast, during the prefill stage, suboptimal data layouts and memory access patterns can substantially reduce computational efficiency.
@@
II Related Works
II-A Light-Weight LLMs
@@
II-B On-device Deployment of LLMs
@@
There are some inference frameworks widely used including llama.cpp [11], MLC LLM [12], mnn-llm [32], mllm [33], etc. Among them,llama.cpp and mnn-llm are mainly optimized for LLM inference on CPUs, while MLC LLM leverages the computational capabilities of GPUs. Despite the availability of dedicated AI accelerators such as NPUs, APUs, and TPUs in the latest generation of mobile SoCs [34], developing for these accelerators presents challenges due to the typically closed-source nature of vendor-specific SDKs. Currently, only mllm and PowerInfer-2 [35] claim to support Qualcomm NPUs for accelerating LLM inference on mobile devices. The development of LLM solutions for AI accelerators with dedicated architectures is still an ongoing challenge for developers.
In addition to those open-source LLM inference engines, mobile SoC vendors and manufacturers have also realized the importance of on-device LLM inference. Qualcomm, the vendor of Snapdragon SoCs, asserts that it can accelerate Llama-2 7B using the NPU on Snapdragon 8 Gen3 [36]. Similarly, MediaTek has announced optimizations for Llama 2 Generative AI applications using its APU on the Dimensity 9300 plus [37]. However, both solutions remain closed-source and have yet to be validated through real-world applications.
II-C Existing Measurement Studies
@@
However, none of these studies provide in-depth analysis about the impact of mobile hardware during inference, such as accelerator utilization, memory bandwidth, and the influence of DVFS on both CPU and GPU, heat buildup during multiple runs, parallel execution with other tasks. Although [9] touches on the impact of SoCs, it only examines the macro differences of SoCs from a single vendor. In contrast, our approach involves testing several SoCs from multiple vendors to gain a comprehensive understanding of how mainstream mobile processors support LLMs. This will help us identify hardware bottlenecks and potential future upgrades necessary for optimizing LLM performance on mobile devices.
@@
| Xiaomi Pad6 Pro | Snapdragon 8+ Gen1 | Cortex-X2/710/510 | Adreno 730 | 12GB/16GB | high-tier | Android13 | 2022.05 |
| Huawei Matepad11 Pro | Snapdragon 870 | Cortex-A77/A77/A55 | Adreno 650 | 5.5GB/8GB | mid-tier | Harmony 4 | 2021.01 |
@@
| Huawei Nova7 | Kirin 985 | Cortex-A76/A76/A55 | Mali-G77 MP8 | 4GB/8GB | mid-tier | Harmony 4 | 2020.04 |
| iPad Pro 11-inch (3rd generation) | Apple M1 | Unknown (8-Core) | Unknown (8-Core) | 5.5GB/8GB | top-tier | iPadOS 18.2.1 | 2021.04 |
III Experimental Setup and Methodology
III-A Experimental Setup
Testing Devices. Table I shows the devices used in our measurement. There are mobile devices from four SoC vendors: Snapdragon, MediaTek, Hisilicon and Apple. All CPUs feature eight Arm Cortex cores but differ in implementation details such as big.LITTLE configurations and cache sizes. It is noteworthy that for non-iOS devices, Snapdragon 8 Gen3 and Dimensity 9300 represent the most advanced SoCs available on mobile devices. We also test devices ranging from high-tier to mid-tier to capture a broad spectrum of hardware heterogeneity. Specifically, the top-tier devices represent the current upper bound for on-device LLM inference; high-tier devices capture the gains from recent hardware improvements, and mid-tier devices test the feasibility of running high-accuracy models on mainstream phones. In terms of GPU, the vast majority of mobile GPUs are from Adreno (Qualcomm), Mali (Arm). Thus, we designate Snapdragon SoCs as representatives of Adreno GPUs and the rest as representatives of Mali GPUs. For iOS device, the iPad selected is equipped with a laptop-level SoC. Because iOS imposes strict per-app memory limits, lower-tier Apple devices cannot feasibly host our target LLM, so we only include a top-tier device.
@@ III-A Experimental Setup
Frameworks and Engines. In this paper, we would like to reveal the performance of LLM on a variety of different accelerators. To achieve this, we select two representative frameworks: llama.cpp for CPU deployments, and MLC LLM for general GPU deployments, to assess the inference capabilities of general accelerators. Additionally, llama.cpp supports GPU for specific SoCs, such as certain Snapdragon SoCs and Apple SoCs. For novel AI accelerators (e.g., NPU), due to the lack of available frameworks, rather than conducting our own device deployments, we refer to existing studies for results of mllm [38] and PowerInfer- [35], which are only available for Qualcomm NPUs. These references will be discussed in subsequent sections.
@@
MLC LLM [12] is a high-performance LLM deployment engine built on the TVM machine learning compiler, MLC LLM is designed to harness the power of mobile GPUs. Leveraging TVM, MLC LLM can automatically generate and optimize operators for a specific model and backend. These operators are then loaded and executed via the TVM Runtime. For Android-based GPUs, MLC LLM utilizes the OpenCL backend to ensure efficient execution.
@@
Measurement toolkits. In order to collect fine-grained hardware metrics such as real-time CPU frequency, utilization, and memory bandwidth on non-Apple devices, we choose Perfetto [39], Snapdragon Profiler [13] and Arm streamline [14] as our primary tools for monitoring hardware dynamics during inference. Perfetto offers comprehensive system-wide performance traces from Android devices, including data from the kernel scheduler, userspace instrumentation, and other sources. For CPU monitoring, Perfetto captures frequency, utilization, and event scheduling for each CPU core. Snapdragon Profiler and Arm Streamline are vendor-specific profilers used to track GPU behavior, including arithmetic and load/store unit utilization. Specifically, Snapdragon Profiler provides detailed timelines of OpenCL kernel executions, which is valuable for operator-level analysis. As for the Apple device, we use Xcode [15] to monitor the memory allocation, real-time bandwidth and GPU utilization.
Evaluation Metrics. We collect results from two main aspects. First, we gather performance metrics that are of interest to users, including prefill and decoding speed, memory footprints, and battery consumption. These results are present in Section IV. Second, we analyze hardware dynamics, focusing on real-time processor utilization and working frequency, impact of Dynamic Voltage and Frequency Scaling (DVFS), heat accumulation and operator implementation, which developers are concerned about. These results are present in Section V. They provide clues for the performance metrics and can help us identify potential hardware and operating system bottlenecks during inference.
III-B Workflow and Methodology
@@
For llama.cpp, we build the model library and compile the source code into an executable file on each device individually, using CMake and the Android NDK. All the compile options are kept the same by default. When necessary, we adjust the compile options to enable specialized instructions to ensure optimal performance (Section V-A2). Since it relies on basic C libraries, no additional environment setup is required. After transferring model weights in GGUF format to the device, inference can be performed easily using the ADB command-line tool. For MLC LLM, there is a native application which TVM runtime and necessary libraries are packed in. The APK can be installed on the device, allowing interaction with the LLM through a graphical interface. MLC LLM cross-compiles the LLM models for the mobile platform, and the runtime version (including tvm and java) is the same on all devices.
@@
IV Performance: Users’ Perspectives
@@
IV-A Token Throughput and Memory Footprints
IV-A1 Performance on CPUs
@@
Figure 3 shows the performance of CPUs, including prefill speed and decoding speed with llama.cpp. Overall, there has been a consistent improvement in both prefill and decoding speeds with new-generation processors. The Apple M1, as a laptop-tier SoC, demonstrates the highest performance. The Dimensity , which ranks first among non-Apple SoCs, shows about a % performance gap compared with the Apple M1 in the decode stage, but delivers only % of the Apple M1’s throughput in the prefill stage. The Snapdragon 8 Gen (Xiaomi Pro) ranks third among all tested devices, showing % throughput of the Dimensity . Comparing Snapdragon SoCs across different tiers, each successive generation shows approximately a improvement in prefill speed, while decode speed improves even more significantly, ranging from to . We note that Snapdragon 870 (Huawei Matepad11 Pro) and Kirin 9000E (Huawei Matepad12.6 Pro) consist of the same cores, and the performance gap may result from higher CPU frequency on Kirin 9000E. Notably, a significant performance gap persists between Kirin SoCs and those from Snapdragon and MediaTek in terms of supporting LLMs.
@@
Our results highlight a notable performance disparity in LLM inference across mobile CPUs. Specifically, top-tier SoCs like the Apple M1 or the Dimensity demonstrate significantly better LLM support compared to mid-tier SoCs. This suggests that for optimal LLM performance, developers should focus on leveraging the most advanced processors available. Furthermore, top-tier SoCs support advanced machine instructions that can significantly boost performance. We will examine this potential in detail in Section V-A2.
@@
We further evaluate the performance with two scenarios: one involving a prompt of tokens with a generation of tokens, and the other involving a prompt of tokens with a generation of tokens. The larger input matrix in the latter case implies greater computational intensity and memory resource demands. Figure 3 illustrates that inference speed decreases as the token length increases. The results may be attributed to the impact of DVFS. With longer prompts and generation tasks, the increased load causes the CPU frequency to throttle more aggressively. Across all tested devices, the performance degradation for both prefill and decoding is relatively consistent, ranging between and . An exception is that the prefill speed achieves a improvement on the Apple M1, suggesting that matrix parallel computation on CPUs still has room for further optimization.
@@
In addition to latency, memory footprint is a also crucial metric. On all six devices, memory usage remains consistently around GB and shows minimal variation during inference once the model weights are loaded into main memory from external storage. This consistency is due to the fact that llama.cpp applies for a fixed-size memory block during initialization, which remains constant once allocation is complete. The allocated memory is then segmented into various components for different variables, including model weights, KV-cache, buffers, context and etc.
Implication 1
LLMs running on mobile CPUs exhibit only slightly additional memory requirements beyond the model weights. Specifically, for a -bit quantized lightweight LLM with B parameters, the maximum memory usage is no more than GB. Moreover, the parameters of many popular lightweight LLMs range from B to B, which further reduces memory requirement. Thus, as long as a mobile device has more than GB available RAM, it is sufficient to run these models locally.
IV-A2 Performance on GPUs